### v1

# Astute RAG: Overcoming Imperfect Retrieval Augmentation and Knowledge Conflicts for Large Language Models

Abstract
Retrieval-Augmented Generation (RAG), while effective in integrating external knowledge to address the limitations of large language models (LLMs), can be undermined by imperfect retrieval, which may introduce irrelevant, misleading, or even malicious information. Despite its importance, previous studies have rarely explored the behavior of RAG through joint analysis on how errors from imperfect retrieval attribute and propagate, and how potential conflicts arise between the LLMs’ internal knowledge and external sources. We find that imperfect retrieval augmentation might be inevitable and quite harmful, through controlled analysis under realistic conditions. We identify the knowledge conflicts between LLM-internal and external knowledge from retrieval as a bottleneck to overcome in the post-retrieval stage of RAG. To render LLMs resilient to imperfect retrieval, we propose Astute RAG, a novel RAG approach that adaptively elicits essential information from LLMs’ internal knowledge, iteratively consolidates internal and external knowledge with source-awareness, and finalizes the answer according to information reliability. Our experiments using Gemini and Claude demonstrate that Astute RAG significantly outperforms previous robustness-enhanced RAG methods. Notably, Astute RAG is the only approach that matches or exceeds the performance of LLMs without RAG under worst-case scenarios. Further analysis reveals that Astute RAG effectively resolves knowledge conflicts, improving the reliability and trustworthiness of RAG systems.
1 Introduction
Retrieval augmented generation (RAG) has become the standard approach for large language models (LLMs) to tackle knowledge-intensive tasks (Guu et al., 2020; Lewis et al., 2020). Prior works mainly leverage RAG to address the inherent knowledge limitations of LLMs, effectively integrating missing information and grounding to reliable sources. However, recent research has highlighted a significant drawback that RAG might rely on imperfect retrieval results, including irrelevant, misleading, or even malicious information, which eventually leads to inaccurate LLM responses (Chen et al., 2024a; Xiang et al., 2024; Zou et al., 2024). For example, when asked about the practice of eating rocks, LLMs might cite misleading information, such as a satirical news source claiming that one should consume at least one rock per day. [^1] https://www.bbc.com/news/articles/cd11gzejgz4o. The occurrence of imperfect retrieval augmentation is inevitable, driven by factors such as corpus quality limitations (Shao et al., 2024), the reliability of retrievers (Dai et al., 2024), and the complexity of the queries (Su et al., 2024). This poses a significant challenge to the trustworthiness of RAG.
While there have been independent analyses of information retrieval and RAG in the context of LLMs (Su et al., 2024; Mallen et al., 2023), previous studies have rarely connected the behaviors of retrieval and subsequent generation, particularly regarding the propagation of information retrieval errors, which may lead to knowledge conflicts (Longpre et al., 2021; Wang et al., 2023a; Xu et al., 2024b) between LLMs and context. To this end, we conduct comprehensive analyses on the occurrence of imperfect retrieval augmentation and its impact on LLM behavior under realistic conditions (Section 2). We conduct controlled experiments on a diverse range of general, domain-specific, and long-tail questions from NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), BioASQ (Tsatsaronis et al., 2015), and PopQA (Mallen et al., 2023). We observe that imperfect retrieval augmentation is widespread even with adept real-world search engine (such as Google Search with Web as corpus) – roughly 70% retrieved passages do not directly contain true answers, leading to the impeded performance of LLM with RAG augmentation. [^2] Note that some passages may contain information indirectly relevant to the answer, but may unintentionally mislead or distract LLMs.
These findings underscore the potential severity of the imperfect retrieval issue in real-world RAG and highlight the widespread existence of knowledge conflicts as the bottleneck to overcome it (Figure 1). Recent studies demonstrate that LLM-internal and external knowledge offer distinct advantages, but LLMs often struggle to consolidate conflicting information reliably, failing to respond based on collective knowledge (Mallen et al., 2023; Tan et al., 2024; Xie et al., 2024; Jin et al., 2024). This raises the following research question: Is there an effective method to combine internal (from LLMs’ pretrained weights) and external (from specific corpora or knowledge bases) knowledge for more reliable RAG? Previous work has widely explored using external knowledge to enhance LLMs through RAG. We seek to further leverage LLMs’ internal knowledge to recover from RAG failures
Figure 1: Knowledge conflicts between the LLMs’ internal knowledge and retrieved knowledge from external sources. We report the overall results with Claude under the setting in Section 4.1.
Motivated by these important real-world challenges, we propose Astute RAG (Section 3), a novel RAG approach designed to be resilient to imperfect retrieval augmentation, while preserving RAG grounding effect when RAG is reliable. To this end, Astute RAG needs effectively differentiate the reliability of the LLM’s intrinsic knowledge and the external information retrieved in RAG, utilizing each only when trustworthy and ensuring proper integration. Specifically, Astute RAG initially elicits information from LLMs’ internal knowledge to explicitly complement the passages retrieved from external sources. Then, Astute RAG conducts source-aware knowledge consolidation of information from various internal and external sources. The desiderata is combining consistent information, identifying conflicting information, and filtering out irrelevant information. Finally, Astute RAG proposes answers based on each group of consistent passages and compares the answers from different passage groups to determine the final answer. Our experiments involving Gemini and Claude [^3] https://www.anthropic.com/claude on various datasets (Section 4) demonstrate the superior performance of Astute RAG compared to previous RAG approaches designed to be robust against retrieval corruptions. Moreover, Astute RAG consistently outperforms baselines across different retrieval quality levels. Notably, Astute RAG is the only RAG method that achieves performance comparable to or even surpassing conventional use of LLMs under the worst-case scenario where all retrieved passages are unhelpful. Further analysis reveals the effectiveness of Astute RAG in resolving knowledge conflicts between internal and external knowledge.
To conclude, our core contributions are threefold. First, we analyze RAG under realistic conditions, identifying imperfect retrieval augmentation as a significant contributor to RAG failures and pinpointing knowledge conflicts as the primary bottleneck in overcoming it. Second, we propose Astute RAG, which explicitly addresses conflicts between LLM-internal and external knowledge, thereby recovering from RAG failures. Third, experiments with various LLMs and datasets demonstrate the effectiveness of Astute RAG, even in the most challenging scenarios.
2 Imperfect Retrieval: The Pitfall of RAG
Figure 2: Imperfect retrieval (samples with low retrieval precision) is prevalent in real-world RAG.
To better showcase the common real-world challenges and to make better motivate for improved methodological designs, we evaluate retrieval quality, end-to-end RAG performance, and knowledge conflicts on a controlled set of data. The selected data encompass a diverse range of general, domain-specific, and long-tail questions from NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), BioASQ (Tsatsaronis et al., 2015), and PopQA (Mallen et al., 2023). Our analysis is based on realistic retrieval results with Google Search [^4] https://developers.google.com/custom-search/v1/overview as the retriever and the Web as the corpus. This setting allows us to analyze the severity of imperfect retrieval in real-world RAG. Overall, we sample 1K short-form QA instances from these datasets, and pair each instance with 10 retrieved passages.
Imperfect retrieval is common. We examine the occurrence of correct answers in retrieved passages as an approximation of retrieval quality. Since we mainly focus on short-form QA which provides most variants of the correct answer for each question, the approximation through string matching can give us a rouge intuition of how precise the retrieval result is. Specifically, we define the retrieval precision as the ratio of passages containing the correct answer for each instance:
Retrieval Precision={number of retrieved passages containing correct answer}{number of total retrieved $\text{Retrieval Precision}=\frac{\text{\{number of retrieved passages containing correct answer\}}}{\text{\{number of total retrieved passages\}}}$
As shown in Figure 2, although instances from different datasets exhibit different data distributions, imperfect retrieval is prevalent. Specifically, $\sim$20of the overall data have no mentions of the correct answer within any retrieved passage, including 34% on NQ, 18% on TriviaQA, 24% on BioASQ, and 50% on PopQA. This finding also aligns with previous observation on information retrieval (Thakur et al., 2024), that highlights that the number of positive passages can be very limited.
Imperfect retrieval leads to RAG failures. We further analyze the relation between retrieval quality and RAG performance. We compare the performance of Claude 3.5 Sonnet, with and without RAG and report the results by retrieval precision in Figure 4. In general, RAG is helpful when the retrieval precision is not lower than 20%. When the retrieval precision is close to 0, the model with RAG performs much worse than without RAG, indicating that imperfect retrieval augmentation can be the cause of RAG failures. This finding aligns with the previous observation from Yu et al. (2024) that adding more retrieved passages does not necessarily lead to better performance, as the additional passages might reduce the retrieval precision.
Knowledge conflicts widely exist in RAG failures. We provide an in-depth analyses of knowledge conflicts between LLMs’ internal knowledge and retrieved passages from external sources. With Claude 3.5 Sonnet as the LLM, Figure 1 shows that 19.2% of the overall data exhibit knowledge conflicts, where either the answer with or without RAG is correct. Among the conflicting cases, the internal knowledge is correct on 47.4% of them, while the external knowledge is correct on the remaining 52.6%. These results emphasize the importance of effectively combining the internal and external knowledge to overcome the inherent limitation of relying solely on either source. However, previous work (Tan et al., 2024; Xie et al., 2024; Jin et al., 2024) show that LLMs might respond based on misleading information rather than comprehensive understanding of the conflicting knowledge in this context.
3 Astute RAG: Overcoming the Pitfall
Figure 3: Overview of the proposed Astute RAG framework. Astute RAG is designed to better combine the information from the external sources (e.g. web, domain-specific corpora, knowledge bases) and internal knowledge of the LLMs by employing a consolidation mechanism to address the information conflicts, which eventually leads to better quality generated outputs.
We begin with formulating the problem of imperfect retrieval in RAG (Section 3.1). We then provide an overview of Astute RAG, designed to overcome this problem (Section 3.2). Subsequently, we delve into the three major steps of Astute RAG, including adaptive generation of internal knowledge (Section 3.3), source-aware knowledge consolidation (Section 3.4), and answer finalization (Section 3.5).
3.1 Problem Formulation
Our objective is to mitigate the effects of imperfect retrieval augmentation, resolve knowledge conflicts between the LLM’s internal knowledge and external sources (such as custom/public corpora and knowledge bases), and ultimately produce more accurate and reliable responses from LLMs.
Given a set of retrieved passages from external sources $E=[e_{1},\ldots,e_{n}]$, a pre-trained LLM $\mathcal{M}$ (accessible through prediction-only APIs, encompassing commercial black-box ones), and a query $q$, the task is to generate the corresponding correct answer $a^{*}$. Notably, this setting is orthogonal to prior work on improving the retriever, training LLMs, or conducting adaptive retrieval, which are mainly preliminary steps.
3.2 Overview of the Framework
Astute RAG is designed to better leverage collective knowledge from both internal knowledge of LLMs and external corpus, for more reliable responses. As shown in Figure 3 and Algorithm 1, Astute RAG starts from acquiring the most accurate, relevant, and thorough passage set from the LLMs’ internal knowledge. Then, internal and external knowledge are consolidated in an iterative way, by comparing the generated and retrieved passages. Finally, the reliability of conflicting information is compared and the final output is generated according to the most reliable knowledge.
Algorithm 1 Astute RAG
1:Query $q$, Retrieved Passages $E=[e_{1},\ldots,e_{n}]$, Large Language Model $\mathcal{M}$, Number of Iteration $t$, Max Number of Generated Passages $\hat{m}$, Prompt Templates $p_{gen},p_{con},p_{ans}$
2:Adaptively generate passages: $I\leftarrow\mathcal{M}(p_{gen},q,\hat{m})$triangleright Section 3.3
3:Combine internal and external passages: $D_{0}\leftarrow E\oplus I$
4:Assign passage sources: $S_{0}\leftarrow[\mathbbm{1}_{\{d\in E\}}\text{for}\ d\ \text{in}\ D_{0}]$
5:if $t>1$ then
6:     for $j=1,\ldots,t-1$ do $\triangleright$ Section 3.4
7:         Consolidate knowledge: $\langle D_{j+1},S_{j+1}\rangle\leftarrow\mathcal{M}(p_{con},q,\langle D_{0},S_{0}\rangle,\langle D_{j},S_{j}\rangle)$
8:     end for
9:     Finally consolidate and answer: $a\leftarrow\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle,\langle D_{t-1},S_{t-1}\rangle)$triangleright Section 3.5
10:else
11:     Consolidate knowledge and finalize the answer: $a\leftarrow\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle)$
12:end if
13:return $a$
3.3 Adaptive Generation of Internal Knowledge
In the first step, we elicit internal knowledge from LLMs. This LLM-internal knowledge, reflecting the consensus from extensive pre-training and instruction-tuning data, can supplement any missing information from the limited set of retrieved passages and enable mutual confirmation between LLM-internal and external knowledge. This is especially valuable when the majority of retrieved passages might be irrelevant or misleading. Specifically, we prompt LLMs to generate passages based on the given question $q$, following Yu et al. (2023a). While Yu et al. (2023a) primarily focused on generating diverse internal passages, we emphasize the importance of reliability and trustworthiness of generated passages. To achieve this goal, we enhance the original method with constitutional principles and adaptive generation.
Inspired by Constitutional AI (Bai et al., 2022), we provide constitutional principles indicating the desired properties of internal passages in the prompt $p_{gen}$ (see Appendix A for details) to guide their generation, emphasizing that the generated passages should be accurate, relevant, and hallucination-free. Moreover, we allow the LLM to perform adaptive generation of passages in its internal knowledge. The LLM can decide how many passages to generate by itself. Rather generating a fix number of passages, we request the LLM to generate at most $\hat{m}$ passages, each covering distinct information, and to directly indicate if no more reliable information is available. This adaptive approach allows the LLM to generate fewer passages (or even no passages at all) when the useful information within internal knowledge is limited and more passages when there are multiple feasible answers in the internal knowledge. In this step, the LLM generates $m\leq\hat{m}$ passages based on its internal knowledge:
$I=[i_{1},\ldots i_{m}]=\mathcal{M}(p_{gen},q,\hat{m}).$
3.4 Iterative Source-aware Knowledge Consolidation
In the second step, we employ the LLM to explicitly consolidate information from both passages generated from its internal knowledge and passages retrieved from external sources. Initially, we combine passages from both internal and external knowledge sources $D_{0}=E\oplus I.$
We additionally ensure source-awareness by providing the source of each passage to LLMs when consolidating knowledge. The source information (internal or external, such as a website) is helpful in assessing the reliability of passages. Here, we provide the passage source as $S_{0}=[\mathbbm{1}_{\{d\in E\}}\text{for}\ d\ \text{in}\ D_{0}].$
To consolidate knowledge, we prompt the LLM (with $p_{con}$ in Appendix A) to identify consistent information across passages, detect conflicting information between each group of consistent passages, and filter out irrelevant information. This step would regroup the unreliable knowledge in input passages into fewer refined passages. The regrouped passages will also attribute their source to the corresponding one or more input passages
$\langle D_{j+1},S_{j+1}\rangle=\mathcal{M}(p_{con},q,\langle D_{0},S_{0}\rangle,\langle D_{j},S_{j}\rangle).$
We find that this is especially helpful in comparing the reliability of conflicting knowledge and addressing knowledge conflicts. Moreover, this knowledge consolidation process can run iteratively for $t$ times to improve the context to be more and more useful. Users can assign a larger number of iterations when the context is lengthy.
3.5 Answer Finalization
In the last step, we prompt the LLM (with $p_{ans}$ in Appendix A) to generate one answer based on each group of passages ($\langle D_{t},S_{t}\rangle$), and then compare their reliability and select the most reliable one as the final answer. This comparison allows the LLM to comprehensively consider knowledge source, cross-source confirmation, frequency, and information thoroughness when making the final decision. Notably, this step can be merged into the last knowledge consolidation step to reduce the inference complexity (the amount of prediction API calls) using a combined prompt:
$a=\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle,\langle D_{t},S_{t}\rangle).$
When $t=1$, the initial passages will be fed into the model directly for knowledge consolidation and subsequent answering: $a=\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle).$
4 Experiments
We evaluate the effectiveness of Astute RAG on overcoming imperfect retrieval augmentation and addressing knowledge conflicts. In this section, we first introduce the experiment setting in detail (Section 4.1). Then, we compare the performance of Astute RAG with various baselines on diverse datasets (Section 4.2). Finally, we provide in-depth analyses (Section 4.3).
Method #API Calls NQ TriviaQA BioASQ PopQA Overall
Claude 3.5 Sonnet (20240620)
No RAG 1 47.12 81.98 50.35 29.78 54.51
RAG 1 44.41 76.68 58.04 35.96 55.47
USC (Chen et al., 2024b) 4 48.14 80.21 61.54 37.64 58.73
GenRead (Yu et al., 2023a) 2 42.03 74.20 56.99 34.27 53.55
RobustRAG (Xiang et al., 2024) 11 47.80 78.09 56.29 37.08 56.53
InstructRAG (Wei et al., 2024) 1 47.12 83.04 58.04 41.01 58.83
Self-Route (Xu et al., 2024a) 1-2 47.46 78.80 59.09 41.01 58.06
Astute RAG (t=1) 2 52.20 84.10 60.14 44.38 61.71
Astute RAG (t=2) 3 53.22 84.45 61.89 44.94 62.67
Astute RAG (t=3) 4 53.56 84.45 62.24 44.94 62.86
Table 1: Main results on Claude under zero-shot setting, showing the accuracy of different benchmark methods vs. Astute RAG, along with their prediction complexity, in number of prediction API calls. Best scores are in bold.
4.1 Experimental Settings
Datasets and metrics. We conduct experiments on the data collected in Section 2 consisting of data from NQ, TriviaQA, BioASQ, and PopQA. For each instance from these datasets, we provide 10 passages collected under a realistic retrieval setting: for each question in our benchmark, we query Google Search to retrieve the top 30 results and select the first 10 accessible websites. From each retrieved website, we extract the paragraph corresponding to the snippet provided in Google Search results as the retrieved passage.. Most of the retrieval results contains natural noise with irrelevant or misleading information. We do not consider enhancements to the retrieval side, such as query rewriting, as such enhancements are typically already incorporated into commercial information retrieval systems. Notably, we do not select questions or annotate answers based on the retrieval results. This setting allows us to analyze the severity of imperfect retrieval in real-world RAG. It distinguishes our benchmark from previous ones that employ synthetic retrieval corruptions or that unintentionally reduce the frequency of imperfect retrieval with biased construction protocols (Chen et al., 2024a; Yang et al., 2024). We also evaluate our method on RGB (Chen et al., 2024a), a RAG diagnostic benchmark evaluating several crucial RAG abilities. Specifically, we choose the English subset of RGB focusing on noise robustness. The benchmark have positive and negative passage sets for each question. We select five negative documents per question as the context to form a worst-case scenario. All the data in these datasets are short-form QA. Following previous work (Xiang et al., 2024; Wei et al., 2024; Mallen et al., 2023), a model response is considered correct if it contains the ground-truth answer. To enhance evaluation reliability, we prompt LLMs to enclose the exact answer within special tokens, extracting them as the final responses.
General Settings of LLMs and RAG. We conduct experiments on two advanced LLMs, including Gemini 1.5 Pro [^5] https://deepmind.google/technologies/gemini/pro/ (gemini-1.5-pro-002) and Claude 3.5 Sonnet [^6] https://www.anthropic.com/news/claude-3-5-sonnet (claude-3-5-sonnet@20240620). The generation temperature is set to 0 and the maximum output tokens is set to 1,024, if not specified otherwise. By default, the passages are presented in the prompt by reversed order. All experiments are under the zero-shot setting for controlled evaluation, where no demonstrations for QA or method-specific steps are provided.
Baselines. We compare Astute RAG with various RAG methods designed for enhanced robustness and representative inference strategies designed to improve response trustworthiness. USC (Chen et al., 2024b) is the universal self-consistency method that samples multiple LLM responses given the same context and aggregates the answers. It provides a reference of naive improvements using additional API calls. The temperature for sampling responses in this baseline is set to 0.7. Genread (Yu et al., 2023a) augments retrieved passages with LLM-generated passages. It provide a reference of presenting passages from both internal and external knowledge in the prompt without effectively combining them. RobustRAG (Xiang et al., 2024) aggregates answers from each independent passage to provide certifiable robustness. We use the keyword aggregation variant as it is shown to be the best-performing variant on advanced LLMs. InstructRAG (Wei et al., 2024) instructs the LLM to provide a rationale connecting the answer with information in passages. For a fair comparison, we use the instructions without training or in-context learning. Self-Route (Xu et al., 2024a) adaptively switches between LLMs with and without RAG. [^7] The original Self-Route switches between RAG and long-context LLMs, while our implementation switches between LLMs with and without RAG to better align with the problem formulation in this paper. This baseline provides a reference of switching between LLMs’ internal and external knowledge.
Implementation Details of Astute RAG. The prompt templates for Astute RAG can be found in Appendix A. By default, we use 2 API calls per query, setting $t=1$ to merge the prompt for knowledge consolidation and answer finalization. For adaptive generation of internal knowledge, we prompt the LLM to generate no more than one passage.
Method #API Calls NQ TriviaQA BioASQ PopQA Overall
Gemini 1.5 Pro (002)
No RAG 1 44.75 80.21 45.80 25.28 51.34
RAG 1 42.71 75.97 55.24 33.71 53.65
USC (Chen et al., 2024b) 4 46.44 76.68 58.39 37.64 56.43
GenRead (Yu et al., 2023a) 2 45.08 77.39 54.90 34.27 54.70
RobustRAG (Xiang et al., 2024) [^8] We observe a high refusal rate in responses of RobustRAG. 11 34.24 67.49 44.06 32.02 45.59
InstructRAG (Wei et al., 2024) 1 46.78 80.57 54.90 34.83 56.14
Self-Route (Xu et al., 2024a) 1-2 47.46 79.86 58.04 38.20 57.58
Astute RAG (t=1) 2 50.17 81.63 58.04 40.45 59.21
Astute RAG (t=2) 3 51.53 81.27 58.74 40.45 59.69
Astute RAG (t=3) 4 48.47 80.21 60.14 42.13 59.21
Table 2: Main results on Gemini under zero-shot setting, showing the accuracy of different benchmark methods vs. Astute RAG, along with their prediction complexity, in number of prediction API calls. Best scores are in bold.
4.2 Main Results
Table 1 and Table 2 presents the results on data with realistic retrieval augmentation for each dataset. By comparing RAG and No RAG, we find that retrieved passages might not always bring benefits to downstream performance – on NQ and TriviaQA, RAG performance lags behind No RAG. We attribute this to that the questions being covered by the LLM’s internal knowledge and the noise in retrieval results misleading the LLM. In contrast, on BioASQ and PopQA, which focus on domain-specific and long-tail questions, RAG significantly improves LLM performance. However, due to imperfect retrieval augmentation, the absolute performance still remains to be unsatisfactory. Among all baselines, no single method consistently outperforms others across all datasets. This observation highlights that these baselines are tailored to distinct settings and may not be universally applicable. For instance, InstructRAG is more effective on TriviaQA, achieving the best performance among all baselines with both Claude and Gemini. In contrast, Self-Route performs better than InstructRAG on both NQ and BioASQ. Moreover, RobustRAG achieves very different performance when applied to Gemini and Claude. Through in-depth analysis, we find that RobustRAG with Gemini exhibits a high refusal rate (refuse to answer) in responses. We attribute this instability to the varying method designs of the baselines, which are tailored for different scenarios, resulting in inconsistent improvement across datasets. Overall, InstructRAG and Self-Route demonstrates the best performance among all baselines when applied to Claude and Gemini respectively. We also note that increasing the number of API calls does not necessarily correlate with improved performance.
Astute RAG consistently outperforms baselines across all datasets of different properties. The overall improvement compared with the best baseline is relatively 6.85% on Claude and 4.13% on Gemini, and the improvements in domain-specific questions are much higher. These results highlight the effectiveness of Astute RAG in overcoming imperfect retrieval augmentation. On Claude, adding more iteration of knowledge consolidation leads to consist improvement. The improvement margin becomes lower when $t$ becomes larger. This is because after each iteration, the remaining improvement space for knowledge consolidation becomes smaller. On Gemini, increasing $t$ primarily benefits BioASQ and PopQA. These two datasets rely more heavily on external knowledge, and iterative knowledge consolidation helps mitigate noise within this external information. Performance on NQ and TriviaQA does not improve further when $t$ reaches 3. We attribute this to the less critical role of external knowledge in these datasets. For setting consistency and efficiency, we set the parameter $\hat{m}$ to a smaller value, limiting the influence of internal knowledge.
4.3 Analyses
Performance by retrieval precision. We compare the performance of Astute RAG and baselines across different subsets partitioned by their retrieval precision, on our collected data with Claude as the LLM. As shown in Figure 4, Astute RAG achieves consistently better performance than all baselines across different retrieval precision, indicating its effectiveness in improving RAG trustworthiness in broad scenarios. Notably, Astute RAG does not sacrifice performance gain under high retrieval quality in exchange for improvement under low retrieval quality. When the retrieval quality is extremely low (close to zero retrieval precision), all other RAG variants underperforms the ’No RAG’ baseline, except for the proposed Astute RAG. This observation aligns with the worst-case results on RGB. It demonstrates the difficulty in overcoming imperfect retrieval augmentation, and verify the effectiveness of Astute RAG in doing so.
(a) Performance by Retrieval Precision
(b) Performance by Knowledge Conflicts
(c) Worst-case Performance on RGB
Figure 4: Analysis results. (a) Performance of Claude across different buckets ranked by retrieval precision. (b) Performance of Claude on conflicting and consistent instances between No RAG and RAG. (c) Worst-case performance of Claude on RGB, where all retrieved passages are negative. Astute RAG reaches a performance close to No RAG, while other RAG systems are far behind.
Effectiveness in addressing knowledge conflicts. We split our collected data in to three subset according to the answers from Claude, with and without RAG. The answers from two inference methods can be both correct, both incorrect, or conflicting with one being correct. These three subsets represents the three situations between internal and external knowledge. The results are shown in Figure 4. On the conflicting subset, Astute RAG successfully chooses the correct answer in approximately 80% of cases, being the most effective method in addressing knowledge conflicts. Notably, Astute RAG even brings performance improvement on the subset where neither internal nor external knowledge alone leads to the correct answer. This indicates that Astute RAG can effectively combine partially-correct information from LLM-internal and external knowledge, to achieve the correct answer through collective information across them.
Worst-case performance on RGB.. Figure 4 presents the results under the worst-case setting on RGB where all retrieved documents are negative. It demonstrates the noise robustness of Astute RAG and baseline RAG methods. The performance gap between RAG and No RAG exceeds 50 points, highlighting the detrimental impact of imperfect retrieval results and emphasizing the importance of providing robust safeguards against worst-case scenarios. While the baseline RAG methods outperform the original RAG, they still obviously fall behind No RAG. Astute RAG is the only RAG method that reaches a performance close to No RAG under the worst-case scenario, further supporting its effectiveness in addressing imperfect retrieval augmentation.
Qualitative study. In Figure 5, we present two representative examples showing the intermediate outputs of Astute RAG. In the first example, LLM without RAG generates a wrong answer, while RAG returns a correct answer. Astute RAG successfully identified the incorrect information in its generated passage and an external passage, avoiding confirmation bias Tan et al. (2024). In the second example, LLM alone is correct, while RAG is incorrect due to the noisy retrieval results. Astute RAG detected the correct answer from noisy retrieved information by checking with its internal knowledge.
Figure 5: Qualitative examples. Top: Astute RAG identified the error in internal knowledge (i.e., generated passage) by confirming with external sources. Bottom: Astute RAG detected the correct answer from noisy retrieved information by checking with its internal knowledge. Standard RAG does not provide an answer because the retrieved passages are too noisy.
5 Related Work
Retrieval augmented generation (RAG) seeks to address the inherent knowledge limitation of LLMs with passages retrieved from external sources of information such as private corpora or public knowledge bases (Guu et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022). Given the widespread adoption of RAG in various real-world applications, including risk-sensitive domains, the negative impact of noisy information within retrieved passages has garnered increasing attention (Cuconasu et al., 2024). Recent work has sought to enhance the robustness of RAG systems against noise from various perspectives, including training LLMs with noisy context (Yu et al., 2023b; Yoran et al., 2024; Pan et al., 2024; Fang et al., 2024), training small models to filter out irrelevant passages (Wang et al., 2023c; Xu et al., 2023), passage reranking (Yu et al., 2024; Glass et al., 2022), dynamic and iterative retrieval (Jiang et al., 2023; Asai et al., 2023; Yan et al., 2024), query rewriting (Ma et al., 2023), and speculative drafting (Wang et al., 2024). These studies focus on distinct modules or stages of RAG systems and are orthogonal to our work.
Our work focuses on enhancing RAG robustness at the post-retrieval stage, after retrieved passages have been provided. On this topic, RobustRAG (Xiang et al., 2024) aggregates answers from each independent passage to provide certifiable robustness. InstructRAG (Wei et al., 2024) instructs the LLM to provide a rationale connecting the answer with information in passages. MADRA (Wang et al., 2023b) applies multi-agent debate to select helpful evidence. However, these works do not explicitly incorporate internal knowledge to recover from RAG failures and may therefore collapse when the majority of retrieved passages are negative. In terms of emphasizing internal knowledge of LLMs in RAG, recent work has explored using LLM-generated passage as context (Yu et al., 2023a; Zhang et al., 2023), adaptively switching between LLMs with and without RAG (Xu et al., 2024a; Mallen et al., 2023; Jeong et al., 2024), and combining answers from internal and external knowledge through contrastive decoding (Zhao et al., 2024; Jin et al., 2024). We focus on a black-box setting where no further training is required, directly addressing knowledge conflicts to combine the helpful information from both sides and achieve more reliable answers.
6 Conclusion
Our paper investigates the impact of imperfect retrieval on the performance of RAG systems and identifies knowledge conflicts as a key challenge. To address this, we introduce Astute RAG, a novel approach that leverages the internal knowledge of LLMs and iteratively refines the generated responses by consolidating internal and external knowledge in a source way. Our empirical results demonstrate the effectiveness of Astute RAG in mitigating the negative effects of imperfect retrieval and improving the robustness of RAG systems, particularly in challenging scenarios with unreliable external sources.
Among the limitations, Astute RAG’s effectiveness hinges on the capabilities of advanced LLMs with strong instruction-following and reasoning abilities, hence potentially more limited applicability with less sophisticated LLMs. As an important future direction, extending the experimental setup to include longer outputs would be important, where the challenges of imperfect retrieval and knowledge conflicts may be even more pronounced. Furthermore, a comprehensive analysis of the impact of various context types (Balachandran et al., 2024) would enhance the understanding of the proposed method’s effectiveness. Future work can also extend our method beyond LLMs and RAG, such as addressing knowledge conflicts in multimodal settings (Zhu et al., 2024).
Acknowledgement
We would like to thank Jinsung Yoon for valuable discussions and insights that helped to improve this paper. We would also like to thank all other colleagues from Google Cloud AI Research for their valuable feedback.
[…]
Appendix A Prompt Template for Astute RAG
Prompt for Adaptive Passage Generation ($p_{gen}$) Generate a document that provides accurate and relevant information to answer the given question. If the information is unclear or uncertain, explicitly state ’I don’t know’ to avoid any hallucinations. Question: {question} Document:
Prompt for Iterative Knowledge Consolidation ($p_{con}$) Task: Consolidate information from both your own memorized documents and externally retrieved documents in response to the given question. * For documents that provide consistent information, cluster them together and summarize the key details into a single, concise document. * For documents with conflicting information, separate them into distinct documents, ensuring each captures the unique perspective or data. * Exclude any information irrelevant to the query. For each new document created, clearly indicate: * Whether the source was from memory or an external retrieval. * The original document numbers for transparency. Initial Context: {context} Last Context: {context} Question: {question} New Context:
Prompt for Knowledge Consolidation and Answer Finalization ($p_{ans}$) Task: Answer a given question using the consolidated information from both your own memorized documents and externally retrieved documents. Step 1: Consolidate information * For documents that provide consistent information, cluster them together and summarize the key details into a single, concise document. * For documents with conflicting information, separate them into distinct documents, ensuring each captures the unique perspective or data. * Exclude any information irrelevant to the query. For each new document created, clearly indicate: * Whether the source was from memory or an external retrieval. * The original document numbers for transparency. Step 2: Propose Answers and Assign Confidence For each group of documents, propose a possible answer and assign a confidence score based on the credibility and agreement of the information. Step 3: Select the Final Answer After evaluating all groups, select the most accurate and well-supported answer. Highlight your exact answer within <<<ANSWER>>> your answer <<</ANSWER>>>. Initial Context: {context_init} [Consolidated Context: {context}] # optional Question: {question} Answer:
Appendix B Data Collection
Encompassing a diverse range of natural questions, our benchmark consists of realistic retrieval results with Google Search [^9] https://developers.google.com/custom-search/v1/overview as the retriever and the Web as the corpus. Notably, we do not select questions or annotate answers based on the retrieval results. This setting allows us to analyze the severity of imperfect retrieval in real-world RAG. It distinguishes our benchmark from previous ones that employ synthetic retrieval corruptions or that unintentionally reduce the frequency of imperfect retrieval with biased construction protocols (Chen et al., 2024a; Yang et al., 2024). Overall, our benchmark contains 1,042 short-form question-answer pairs, each paired with 10 retrieved passages.
Question-answer pairs. We consider question-answer pairs from four datasets of different properties spanning across general questions, domain-specific questions, and long-tail questions. NQ (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017) are two widely-studied question-answering (QA) datasets in general domains. BioASQ (Tsatsaronis et al., 2015) is from biomedical domain that has demonstrated significant benefits from RAG when general-purpose LLMs are considered. PopQA (Mallen et al., 2023) focuses on long-tail knowledge and has been shown to be challenging for even advanced LLMs to solve without external knowledge. All these datasets contain questions with short-form answers and most of them list all valid answer variants. This format can support automatic verification of answer appearance in retrieved passages and model responses, leading to more precise evaluations.
Retrieval process. For each question in our benchmark, we query Google Search to retrieve the top 30 results and select the first 10 accessible websites. From each retrieved website, we extract the paragraph corresponding to the snippet provided in Google Search results as the retrieved passage. We do not consider enhancements to the retrieval side, such as query rewriting, as such enhancements are typically already incorporated into commercial information retrieval systems.

### v2

@@
 Abstract
-Retrieval-Augmented Generation (RAG), while effective in integrating external knowledge to address the limitations of large language models (LLMs), can be undermined by imperfect retrieval, which may introduce irrelevant, misleading, or even malicious information.
-Despite its importance, previous studies have rarely explored the behavior of RAG through joint analysis on how errors from imperfect retrieval attribute and propagate, and how potential conflicts arise between the LLMs’ internal knowledge and external sources.
-We find that imperfect retrieval augmentation might be inevitable and quite harmful, through controlled analysis under realistic conditions.
-We identify the knowledge conflicts between LLM-internal and external knowledge from retrieval as a bottleneck to overcome in the post-retrieval stage of RAG.
+Retrieval-augmented generation (RAG), while effective in integrating external knowledge to enhance large language models (LLMs), can be undermined by imperfect retrieval, which may introduce irrelevant, misleading, or even malicious information.
+Despite its importance, previous studies have rarely explored the behavior of RAG with errors from imperfect retrieval, and how potential conflicts arise between the LLMs’ internal knowledge and external sources.
+We show that imperfect retrieval augmentation might be inevitable and quite harmful, through controlled analysis under realistic conditions.
+Knowledge conflicts between LLM-internal and external knowledge from retrieval is a bottleneck to overcome in the post-retrieval stage of RAG.
 To render LLMs resilient to imperfect retrieval, we propose Astute RAG, a novel RAG approach that adaptively elicits essential information from LLMs’ internal knowledge, iteratively consolidates internal and external knowledge with source-awareness, and finalizes the answer according to information reliability.
-Our experiments using Gemini and Claude demonstrate that Astute RAG significantly outperforms previous robustness-enhanced RAG methods.
+Our experiments with Gemini and Claude demonstrate that Astute RAG significantly outperforms previous robustness-enhanced RAG methods.
 Notably, Astute RAG is the only approach that matches or exceeds the performance of LLMs without RAG under worst-case scenarios.
-Further analysis reveals that Astute RAG effectively resolves knowledge conflicts, improving the reliability and trustworthiness of RAG systems.
+Astute RAG effectively resolves knowledge conflicts, improving the reliability and trustworthiness of RAG systems.
+Astute RAG: Overcoming Imperfect Retrieval Augmentation and Knowledge Conflicts for Large Language Models
+Fei Wang12∗  Xingchen Wan1  Ruoxi Sun1  Jiefeng Chen1  Sercan Ö. Arık1 1Google Cloud AI Research  2University of Southern California fwang598@usc.edu{xingchenw,ruoxis,jiefengc,soarik}@google.com
+11footnotetext: Work done during internship at Google.
 1 Introduction
-Retrieval augmented generation (RAG) has become the standard approach for large language models (LLMs) to tackle knowledge-intensive tasks (Guu et al., 2020; Lewis et al., 2020).
+Retrieval-augmented generation (RAG) is commonly used for large language models (LLMs) to tackle knowledge-intensive tasks (Guu et al., 2020; Lewis et al., 2020).
 Prior works mainly leverage RAG to address the inherent knowledge limitations of LLMs, effectively integrating missing information and grounding to reliable sources.
-However, recent research has highlighted a significant drawback that RAG might rely on imperfect retrieval results, including irrelevant, misleading, or even malicious information, which eventually leads to inaccurate LLM responses (Chen et al., 2024a; Xiang et al., 2024; Zou et al., 2024).
-For example, when asked about the practice of eating rocks, LLMs might cite misleading information, such as a satirical news source claiming that one should consume at least one rock per day.
-[^1] https://www.bbc.com/news/articles/cd11gzejgz4o.
-The occurrence of imperfect retrieval augmentation is inevitable, driven by factors such as corpus quality limitations (Shao et al., 2024), the reliability of retrievers (Dai et al., 2024), and the complexity of the queries (Su et al., 2024).
-This poses a significant challenge to the trustworthiness of RAG.
-While there have been independent analyses of information retrieval and RAG in the context of LLMs (Su et al., 2024; Mallen et al., 2023), previous studies have rarely connected the behaviors of retrieval and subsequent generation, particularly regarding the propagation of information retrieval errors, which may lead to knowledge conflicts (Longpre et al., 2021; Wang et al., 2023a; Xu et al., 2024b) between LLMs and context.
-To this end, we conduct comprehensive analyses on the occurrence of imperfect retrieval augmentation and its impact on LLM behavior under realistic conditions (Section 2).
-We conduct controlled experiments on a diverse range of general, domain-specific, and long-tail questions from NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), BioASQ (Tsatsaronis et al., 2015), and PopQA (Mallen et al., 2023).
-We observe that imperfect retrieval augmentation is widespread even with adept real-world search engine (such as Google Search with Web as corpus) – roughly 70% retrieved passages do not directly contain true answers, leading to the impeded performance of LLM with RAG augmentation.
-[^2] Note that some passages may contain information indirectly relevant to the answer, but may unintentionally mislead or distract LLMs.
-These findings underscore the potential severity of the imperfect retrieval issue in real-world RAG and highlight the widespread existence of knowledge conflicts as the bottleneck to overcome it (Figure 1).
-Recent studies demonstrate that LLM-internal and external knowledge offer distinct advantages, but LLMs often struggle to consolidate conflicting information reliably, failing to respond based on collective knowledge (Mallen et al., 2023; Tan et al., 2024; Xie et al., 2024; Jin et al., 2024).
-This raises the following research question: Is there an effective method to combine internal (from LLMs’ pretrained weights) and external (from specific corpora or knowledge bases) knowledge for more reliable RAG?
-Previous work has widely explored using external knowledge to enhance LLMs through RAG.
-We seek to further leverage LLMs’ internal knowledge to recover from RAG failures
-Figure 1: Knowledge conflicts between the LLMs’ internal knowledge and retrieved knowledge from external sources. We report the overall results with Claude under the setting in Section 4.1.
-Motivated by these important real-world challenges, we propose Astute RAG (Section 3), a novel RAG approach designed to be resilient to imperfect retrieval augmentation, while preserving RAG grounding effect when RAG is reliable.
-To this end, Astute RAG needs effectively differentiate the reliability of the LLM’s intrinsic knowledge and the external information retrieved in RAG, utilizing each only when trustworthy and ensuring proper integration.
-Specifically, Astute RAG initially elicits information from LLMs’ internal knowledge to explicitly complement the passages retrieved from external sources.
-Then, Astute RAG conducts source-aware knowledge consolidation of information from various internal and external sources.
+However, recent research has highlighted a significant drawback that RAG might rely on imperfect retrieval, including irrelevant, misleading, or even malicious information (Fig.
+1), which eventually leads to inaccurate LLM responses (Chen et al., 2024a; Zou et al., 2024).
+Moreover, recent studies have shown that retrieval augmentation can confuse LLMs when retrieved passages are conflicting with LLMs’ parametric knowledge Tan et al.
+(2024); Xie et al.
+(2024); Jin et al.
+(2024).
+These pose significant challenges to the trustworthiness of RAG.
+Figure 1: Knowledge conflicts between the LLMs’ internal knowledge and retrieved knowledge from external sources. We report the overall results with Claude under the setting in Sec. 5.1.
+To address imperfect retrieval, earlier work seeks to improve the retrieval approaches, such as dynamic and iterative retrieval Jiang et al.
+(2023); Asai et al.
+(2023); Yan et al.
+(2024).
+However, the occurrence of imperfect retrieval is still inevitable, due to corpus quality limitations (Shao et al., 2024), the reliability of retrievers (Dai et al., 2024), and the complexity of queries (Su et al., 2024).
+Consequently, recent work shifts the focus to the generation stage, seeking to reduce the negative impact of noisy retrieved passages Xiang et al.
+(2024); Wei et al.
+(2024).
+Another line of research at generation stage, motivated by knowledge conflicts, has explored complementing retrieved passages with LLM-generated passages Yu et al.
+(2023a); Zhang et al.
+(2023) or deactivating RAG when the retrieved passages are of insufficient quality Xu et al.
+(2024); Mallen et al.
+(2023); Jeong et al.
+(2024).
+Despite the previous work on the impact of imperfect retrieval and knowledge conflicts at RAG generation stage, quantitative analyses lack on two crucial real-world aspects: (i) the relation between retrieval quality and occurrence of knowledge conflicts, and (ii) the extent to which retrieved passages and LLMs’ parametric knowledge can correct each other.
+Method-wise, existing approaches for mitigating RAG failures caused by imperfect retrieval and knowledge conflicts have not yet yielded a training-free method capable of explicitly analyzing conflicting knowledge across various internal and external sources, and achieving worst-case robustness for black-box LLMs.
+In this paper, we first conduct comprehensive analyses to investigate the relation between imperfect retrieval and knowledge conflicts, and examine the frequency of external and LLMs’ internal knowledge mutually correcting each other (Sec.
+3).
+On a diverse range of general, domain-specific, and long-tail questions from NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), BioASQ (Tsatsaronis et al., 2015), and PopQA (Mallen et al., 2023), we observe that imperfect retrieval is widespread even with an adept real-world search engine, leading to the impeded performance of RAG.111such as Google Search with Web as corpus Retrieval precision is tightly correlated with the knowledge conflict rate.
+Mutual correction between the LLM’s knowledge and external knowledge is crucial for recovering from RAG failures.
+Our findings underscore the potential severity of imperfect retrieval in real-world RAG and highlight the widespread existence of knowledge conflicts as the bottleneck.
+We propose Astute RAG, a novel RAG approach designed for resilience to imperfect retrieval augmentation, while preserving RAG grounding effect when retrieval is reliable (Sec.
+4).
+Astute RAG effectively differentiates between consistent and conflicting information from the LLM’s internal knowledge and the externally retrieved passages, assesses their reliability, and ensures proper integration of trustworthy information.
+Astute RAG first adaptively elicits LLMs’ knowledge and then conducts source-aware knowledge consolidation.
 The desiderata is combining consistent information, identifying conflicting information, and filtering out irrelevant information.
-Finally, Astute RAG proposes answers based on each group of consistent passages and compares the answers from different passage groups to determine the final answer.
-Our experiments involving Gemini and Claude [^3] https://www.anthropic.com/claude on various datasets (Section 4) demonstrate the superior performance of Astute RAG compared to previous RAG approaches designed to be robust against retrieval corruptions.
-Moreover, Astute RAG consistently outperforms baselines across different retrieval quality levels.
-Notably, Astute RAG is the only RAG method that achieves performance comparable to or even surpassing conventional use of LLMs under the worst-case scenario where all retrieved passages are unhelpful.
-Further analysis reveals the effectiveness of Astute RAG in resolving knowledge conflicts between internal and external knowledge.
-To conclude, our core contributions are threefold.
-First, we analyze RAG under realistic conditions, identifying imperfect retrieval augmentation as a significant contributor to RAG failures and pinpointing knowledge conflicts as the primary bottleneck in overcoming it.
-Second, we propose Astute RAG, which explicitly addresses conflicts between LLM-internal and external knowledge, thereby recovering from RAG failures.
-Third, experiments with various LLMs and datasets demonstrate the effectiveness of Astute RAG, even in the most challenging scenarios.
-2 Imperfect Retrieval: The Pitfall of RAG
+Finally, Astute RAG proposes answers based on consistent information and compares them to determine the final answer.
+Our experiments with various LLMs (Claude, Gemini and Mistral), demonstrate superior performance of Astute RAG compared to previous RAG approaches designed for robustness (Sec.
+5).
+Astute RAG consistently outperforms baselines across different retrieval quality levels.
+Notably, Astute RAG is the only RAG method that achieves performance comparable to or even surpassing retrieval-free mode of LLMs under the worst-case scenario where all retrieved passages are unhelpful.
+Further analysis reveals the effectiveness of Astute RAG in resolving knowledge conflicts.
+In summary, our core contributions are threefold.
+First, we provide quantitative analyses and novel insights for the connection among imperfect retrieval, knowledge conflicts, and RAG failures under real-world conditions.
+Second, we propose Astute RAG, which explicitly analyzes LLM-internal and external knowledge in-context, assesses their reliablity, and recovers from RAG failures with black-box access.
+Third, with experiments with various LLMs and datasets, we demonstrate the effectiveness of Astute RAG in improving robustness and trustworthiness, even in the most challenging scenarios.
+2 Related Work
 Figure 2: Imperfect retrieval (samples with low retrieval precision) is prevalent in real-world RAG.
-To better showcase the common real-world challenges and to make better motivate for improved methodological designs, we evaluate retrieval quality, end-to-end RAG performance, and knowledge conflicts on a controlled set of data.
-The selected data encompass a diverse range of general, domain-specific, and long-tail questions from NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), BioASQ (Tsatsaronis et al., 2015), and PopQA (Mallen et al., 2023).
-Our analysis is based on realistic retrieval results with Google Search [^4] https://developers.google.com/custom-search/v1/overview as the retriever and the Web as the corpus.
-This setting allows us to analyze the severity of imperfect retrieval in real-world RAG.
-Overall, we sample 1K short-form QA instances from these datasets, and pair each instance with 10 retrieved passages.
-Imperfect retrieval is common.
-We examine the occurrence of correct answers in retrieved passages as an approximation of retrieval quality.
-Since we mainly focus on short-form QA which provides most variants of the correct answer for each question, the approximation through string matching can give us a rouge intuition of how precise the retrieval result is.
-Specifically, we define the retrieval precision as the ratio of passages containing the correct answer for each instance:
-Retrieval Precision={number of retrieved passages containing correct answer}{number of total retrieved $\text{Retrieval Precision}=\frac{\text{\{number of retrieved passages containing correct answer\}}}{\text{\{number of total retrieved passages\}}}$
-As shown in Figure 2, although instances from different datasets exhibit different data distributions, imperfect retrieval is prevalent.
-Specifically, $\sim$20of the overall data have no mentions of the correct answer within any retrieved passage, including 34% on NQ, 18% on TriviaQA, 24% on BioASQ, and 50% on PopQA.
-This finding also aligns with previous observation on information retrieval (Thakur et al., 2024), that highlights that the number of positive passages can be very limited.
-Imperfect retrieval leads to RAG failures.
-We further analyze the relation between retrieval quality and RAG performance.
-We compare the performance of Claude 3.5 Sonnet, with and without RAG and report the results by retrieval precision in Figure 4.
-In general, RAG is helpful when the retrieval precision is not lower than 20%.
-When the retrieval precision is close to 0, the model with RAG performs much worse than without RAG, indicating that imperfect retrieval augmentation can be the cause of RAG failures.
-This finding aligns with the previous observation from Yu et al.
-(2024) that adding more retrieved passages does not necessarily lead to better performance, as the additional passages might reduce the retrieval precision.
-Knowledge conflicts widely exist in RAG failures.
-We provide an in-depth analyses of knowledge conflicts between LLMs’ internal knowledge and retrieved passages from external sources.
-With Claude 3.5 Sonnet as the LLM, Figure 1 shows that 19.2% of the overall data exhibit knowledge conflicts, where either the answer with or without RAG is correct.
+Figure 3: Conflicting rate between answers from LLMs with and without RAG on different retrieval precision.
+RAG aims to address the inherent knowledge limitation of LLMs with passages retrieved from external sources of information such as private corpora or public knowledge bases (Guu et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022).
+Given the widespread real-world adoption of RAG, including risk-sensitive domains, the negative impact of noisy information within retrieved passages has garnered increasing attention (Cuconasu et al., 2024).
+Recent work explored enhancing the robustness of RAG systems against noise from various perspectives, including training LLMs with noisy context (Yu et al., 2023b; Yoran et al., 2024; Pan et al., 2024; Fang et al., 2024), training small models to filter out irrelevant passages (Wang et al., 2023b; Xu et al., 2023), passage reranking (Yu et al., 2024; Glass et al., 2022), dynamic and iterative retrieval (Jiang et al., 2023; Asai et al., 2023; Yan et al., 2024), query rewriting (Ma et al., 2023), and speculative drafting (Wang et al., 2024).
+These focus on distinct modules or stages of RAG systems and are orthogonal to our work.
+Our work focuses on enhancing RAG robustness at the post-retrieval stage, after retrieved passages have been provided.
+On this, RobustRAG (Xiang et al., 2024) aggregates answers from each independent passage to provide certifiable robustness.
+InstructRAG (Wei et al., 2024) instructs the LLM to provide a rationale connecting the answer with information in passages.
+MADRA (Wang et al., 2023a) applies multi-agent debate to select helpful evidence.
+However, these do not explicitly incorporate internal knowledge to recover from RAG failures and therefore might severely suffer when the majority of retrieved passages have issues.
+For emphasizing internal knowledge of LLMs in RAG, recent work explored using LLM-generated passage as context (Yu et al., 2023a), training models to match generated and retrieved passages (Zhang et al., 2023), adaptively switching between LLMs with and without RAG (Xu et al., 2024; Mallen et al., 2023; Jeong et al., 2024), and combining answers through contrastive decoding (Zhao et al., 2024; Jin et al., 2024).
+Different from prior work, we provide a systematic framework on connecting imperfect retrieval, knowledge conflicts, and RAG failures.
+Specifically focusing on the imperfect context setting, our method is training-free and applicable to black-box LLMs, explicitly analyzes internal and external knowledge in-context, and offers broader usability and adaptability.
+3 The Pitfall of RAG
+To better showcase common real-world challenges and motivate improved methodological designs, we evaluate retrieval quality, the occurrence of knowledge conflicts, their relationship, and the mutual correction between external and internal knowledge using a controlled dataset derived from NQ, TriviaQA, BioASQ, and PopQA, datasets widely used for RAG in prior work Xiang et al.
+(2024); Wei et al.
+(2024); Asai et al.
+(2023).
+Different from prior work, our analysis is based on real-world retrieval results with Google Search [^2] https://developers.google.com/custom-search/v1/overview as the retriever and the Web as the corpus.
+Overall, we sample 1K instances, each with 10 retrieved passages.
+Figure 4: Overview of the Astute RAG framework.
+Astute RAG is designed to better combine the information from the external sources (e.g.
+web, domain-specific corpora) and internal knowledge of the LLMs by employing a consolidation mechanism to address the conflicts, which eventually leads to superior generation quality.
+Imperfect retrieval and knowledge conflicts are common and harmful.
+Our initial observations are consist with prior work.
+As shown in Fig.
+2, the retrieval precision [^3] Ratio of passages directly contain true answers.
+is generally low - roughly 70% retrieved passages do not directly contain true answers, consistent with prior work demonstrating the often imperfect nature of retrieval results Thakur et al.
+(2024); Su et al.
+(2024).
+With Claude 3.5 Sonnet as the LLM, Fig.
+1 shows that 19.2% of the overall data exhibit knowledge conflicts, consistent with prior work demonstrating the prevalence of such conflicts across various scenarios Pham et al.
+(2024); Xie et al.
+(2024); Longpre et al.
+(2021).
+Moreover, we observe strong correlations between retrieval precision and RAG performance (Fig.
+7) and between the occurrence of knowledge conflicts and RAG performance (Fig.
+8), findings consistent with prior work on these respective topics Chen et al.
+(2024a); Xie et al.
+(2024).
+Lower retrieval precision increases knowledge conflicts in general.
+As shown in Fig.
+3, most advanced LLMs exhibit the highest conflict rates when retrieval precision is as low as 10%.
+Subsequently, the conflict rate generally decreases as precision increases, although some fluctuations may occur.
+This trend is generally applicable to the studied LLMs with different training processes.
+Notably, when retrieval precision is 0%, conflict rates tend to be significantly lower.
+This suggests that limited external knowledge for the query results in more irrelevant passages rather than incorrect ones.
+Internal and external knowledge can correct each other to a comparable extent.
 Among the conflicting cases, the internal knowledge is correct on 47.4% of them, while the external knowledge is correct on the remaining 52.6%.
 These results emphasize the importance of effectively combining the internal and external knowledge to overcome the inherent limitation of relying solely on either source.
-However, previous work (Tan et al., 2024; Xie et al., 2024; Jin et al., 2024) show that LLMs might respond based on misleading information rather than comprehensive understanding of the conflicting knowledge in this context.
-3 Astute RAG: Overcoming the Pitfall
-Figure 3: Overview of the proposed Astute RAG framework.
-Astute RAG is designed to better combine the information from the external sources (e.g.
-web, domain-specific corpora, knowledge bases) and internal knowledge of the LLMs by employing a consolidation mechanism to address the information conflicts, which eventually leads to better quality generated outputs.
-We begin with formulating the problem of imperfect retrieval in RAG (Section 3.1).
-We then provide an overview of Astute RAG, designed to overcome this problem (Section 3.2).
-Subsequently, we delve into the three major steps of Astute RAG, including adaptive generation of internal knowledge (Section 3.3), source-aware knowledge consolidation (Section 3.4), and answer finalization (Section 3.5).
-3.1 Problem Formulation
-Our objective is to mitigate the effects of imperfect retrieval augmentation, resolve knowledge conflicts between the LLM’s internal knowledge and external sources (such as custom/public corpora and knowledge bases), and ultimately produce more accurate and reliable responses from LLMs.
-Given a set of retrieved passages from external sources $E=[e_{1},\ldots,e_{n}]$, a pre-trained LLM $\mathcal{M}$ (accessible through prediction-only APIs, encompassing commercial black-box ones), and a query $q$, the task is to generate the corresponding correct answer $a^{*}$.
-Notably, this setting is orthogonal to prior work on improving the retriever, training LLMs, or conducting adaptive retrieval, which are mainly preliminary steps.
-3.2 Overview of the Framework
-Astute RAG is designed to better leverage collective knowledge from both internal knowledge of LLMs and external corpus, for more reliable responses.
-As shown in Figure 3 and Algorithm 1, Astute RAG starts from acquiring the most accurate, relevant, and thorough passage set from the LLMs’ internal knowledge.
-Then, internal and external knowledge are consolidated in an iterative way, by comparing the generated and retrieved passages.
-Finally, the reliability of conflicting information is compared and the final output is generated according to the most reliable knowledge.
+However, previous work (Tan et al., 2024; Xie et al., 2024; Jin et al., 2024) shows that LLMs often select knowledge based on unreliable shortcuts, so simply presenting LLM-generated passages in the context may not help.
+4 Astute RAG
 Algorithm 1 Astute RAG
 1:Query $q$, Retrieved Passages $E=[e_{1},\ldots,e_{n}]$, Large Language Model $\mathcal{M}$, Number of Iteration $t$, Max Number of Generated Passages $\hat{m}$, Prompt Templates $p_{gen},p_{con},p_{ans}$
-2:Adaptively generate passages: $I\leftarrow\mathcal{M}(p_{gen},q,\hat{m})$triangleright Section 3.3
+2:Adaptively generate passages: $I\leftarrow\mathcal{M}(p_{gen},q,\hat{m})$triangleright Sec. 4.2
 3:Combine internal and external passages: $D_{0}\leftarrow E\oplus I$
@@
 5:if $t>1$ then
-6:     for $j=1,\ldots,t-1$ do $\triangleright$ Section 3.4
+6:     for $j=1,\ldots,t-1$ do $\triangleright$ Sec. 4.3
 7:         Consolidate knowledge: $\langle D_{j+1},S_{j+1}\rangle\leftarrow\mathcal{M}(p_{con},q,\langle D_{0},S_{0}\rangle,\langle D_{j},S_{j}\rangle)$
 8:     end for
-9:     Finally consolidate and answer: $a\leftarrow\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle,\langle D_{t-1},S_{t-1}\rangle)$triangleright Section 3.5
+9:     Finally consolidate and answer: $a\leftarrow\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle,\langle D_{t-1},S_{t-1}\rangle)$triangleright Sec. 4.4
 10:else
@@
 13:return $a$
+We first provide an overview of Astute RAG (Sec.
+4.1).
+Subsequently, we delve into the three major steps of Astute RAG, including adaptive generation of internal knowledge (Sec.
+4.2), source-aware knowledge consolidation (Sec.
+4.3), and answer finalization (Sec.
+4.4).
+4.1 Overview
+Our objective is to mitigate the effects of imperfect retrieval augmentation, resolve knowledge conflicts between the LLM’s internal knowledge and external sources (such as custom/public corpora and knowledge bases), and ultimately produce more accurate and reliable responses from LLMs.
+Given a set of retrieved passages from external sources $E=[e_{1},\ldots,e_{n}]$, a pre-trained LLM $\mathcal{M}$ (accessible through prediction-only APIs, encompassing commercial black-box ones), and a query $q$, the task is to generate the corresponding correct answer $a^{*}$.
+Notably, this setting is orthogonal to prior work on improving the retriever, training LLMs, or conducting adaptive retrieval, which are mainly preliminary steps.
+Astute RAG is designed to better leverage collective knowledge from both internal knowledge of LLMs and external corpus, for more reliable responses.
+As shown in Fig.
+4 and Alg.
+1, Astute RAG starts from acquiring the most accurate, relevant, and thorough passage set from the LLMs’ internal knowledge.
+Then, internal and external knowledge are consolidated in an iterative way, by comparing the generated and retrieved passages.
+Finally, the reliability of conflicting information is compared and the final output is generated according to the most reliable knowledge.
 4.2 Adaptive Generation of Internal Knowledge
@@
 To achieve this goal, we enhance the original method with constitutional principles and adaptive generation.
-Inspired by Constitutional AI (Bai et al., 2022), we provide constitutional principles indicating the desired properties of internal passages in the prompt $p_{gen}$ (see Appendix A for details) to guide their generation, emphasizing that the generated passages should be accurate, relevant, and hallucination-free.
+Inspired by Bai et al.
+(2022), we provide constitutional principles indicating the desired properties of internal passages in the prompt $p_{gen}$ (see Appx.
+A for details) to guide their generation, emphasizing that the generated passages should be accurate, relevant, and hallucination-free.
 Moreover, we allow the LLM to perform adaptive generation of passages in its internal knowledge.
@@
 The source information (internal or external, such as a website) is helpful in assessing the reliability of passages.
-Here, we provide the passage source as $S_{0}=[\mathbbm{1}_{\{d\in E\}}\text{for}\ d\ \text{in}\ D_{0}].$
-To consolidate knowledge, we prompt the LLM (with $p_{con}$ in Appendix A) to identify consistent information across passages, detect conflicting information between each group of consistent passages, and filter out irrelevant information.
+Here, we provide the passage source as $S_{0}=[\mathbbm{1}_{\{d\in E\}}\text{for}\ d\ \text{in}\ D_{0}].$ To consolidate knowledge, we prompt the LLM (with $p_{con}$ in Appx.
+A) to identify consistent information across passages, detect conflicting information between each group of consistent passages, and filter out irrelevant information.
 This step would regroup the unreliable knowledge in input passages into fewer refined passages.
-The regrouped passages will also attribute their source to the corresponding one or more input passages
+The regrouped passages also attribute their source to the corresponding input passages:
 $\langle D_{j+1},S_{j+1}\rangle=\mathcal{M}(p_{con},q,\langle D_{0},S_{0}\rangle,\langle D_{j},S_{j}\rangle).$
 We find that this is especially helpful in comparing the reliability of conflicting knowledge and addressing knowledge conflicts.
-Moreover, this knowledge consolidation process can run iteratively for $t$ times to improve the context to be more and more useful.
-Users can assign a larger number of iterations when the context is lengthy.
+This knowledge consolidation process can run iteratively for $t$ times to improve better utilization of the retrieved context.
 4.4 Answer Finalization
-In the last step, we prompt the LLM (with $p_{ans}$ in Appendix A) to generate one answer based on each group of passages ($\langle D_{t},S_{t}\rangle$), and then compare their reliability and select the most reliable one as the final answer.
+In the last step, we prompt the LLM (with $p_{ans}$ in Appx.
+A) to generate one answer based on each group of passages ($\langle D_{t},S_{t}\rangle$), and then compare their reliability and select the most reliable one as the final answer.
 This comparison allows the LLM to comprehensively consider knowledge source, cross-source confirmation, frequency, and information thoroughness when making the final decision.
@@
 $a=\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle,\langle D_{t},S_{t}\rangle).$
-When $t=1$, the initial passages will be fed into the model directly for knowledge consolidation and subsequent answering: $a=\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle).$
+When $t=1$, the initial passages will be input to the model directly for knowledge consolidation and subsequent answering: $a=\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle).$
 5 Experiments
+Method NQ TriviaQA BioASQ PopQA Overall NQ TriviaQA BioASQ PopQA Overall
+Claude 3.5 Sonnet (20240620) Gemini 1.5 Pro (002)
+No RAG 47.1 82.0 50.4 29.8 54.5 44.8 80.2 45.8 25.3 51.3
+RAG 44.4 76.7 58.0 36.0 55.5 42.7 76.0 55.2 33.7 53.7
+USC (Chen et al., 2024b) 48.1 80.2 61.5 37.6 58.7 46.4 76.7 58.4 37.6 56.4
+GenRead (Yu et al., 2023a) 42.0 74.2 57.0 34.3 53.6 45.1 77.4 54.9 34.3 54.7
+RobustRAG (Xiang et al., 2024) 47.8 78.1 56.3 37.1 56.5 34.2 67.5 44.1 32.0 45.6444We observe a high refusal rate in RobustRAG for Gemini.
+InstructRAG (Wei et al., 2024) 47.1 83.0 58.0 41.0 58.8 46.8 80.6 54.9 34.8 56.1
+Self-Route (Xu et al., 2024) 47.5 78.8 59.1 41.0 58.1 47.5 79.9 58.0 38.2 57.6
+Astute RAG 52.2 84.1 60.1 44.4 61.7 50.2 81.6 58.0 40.5 59.2
+Mistral-Large (2407), 128B Mistral-Nemo (2407), 12B
+No RAG 46.8 79.5 43.7 24.7 51.1 29.8 67.8 34.3 23.0 40.2
+RAG 43.1 77.4 55.9 36.0 54.7 39.3 66.8 49.0 32.6 48.3
+USC (Chen et al., 2024b) 51.2 80.9 61.5 36.0 59.5 29.5 66.1 36.0 20.2 39.6
+GenRead (Yu et al., 2023a) 40.7 73.1 55.6 35.4 52.7 38.6 68.9 48.3 33.7 48.7
+RobustRAG (Xiang et al., 2024) 42.7 77.7 50.4 34.8 53.0 35.6 71.7 44.1 27.5 46.4
+InstructRAG (Wei et al., 2024) 45.4 80.6 57.3 36.5 56.7 38.3 61.8 50.4 23.6 45.5
+Self-Route (Xu et al., 2024) 45.4 77.7 57.3 38.2 56.2 41.4 73.5 51.8 30.9 51.2
+Astute RAG 50.2 82.7 58.4 42.1 59.9 42.7 73.9 49.3 32.6 51.3
+Table 1: Main results on Claude 3.5 Sonnet, Gemini 1.5 Pro, Mistral-Large, and Mistral-Nemo under zero-shot setting, showing the accuracy of benchmarked alternatives vs.
+Astute RAG.
+Best scores are in bold.
+Note that USC consumes approximately three times more tokens than other RAG methods, and is not directly comparable.
 We evaluate the effectiveness of Astute RAG on overcoming imperfect retrieval augmentation and addressing knowledge conflicts.
-In this section, we first introduce the experiment setting in detail (Section 4.1).
-Then, we compare the performance of Astute RAG with various baselines on diverse datasets (Section 4.2).
-Finally, we provide in-depth analyses (Section 4.3).
-Method #API Calls NQ TriviaQA BioASQ PopQA Overall
-Claude 3.5 Sonnet (20240620)
-No RAG 1 47.12 81.98 50.35 29.78 54.51
-RAG 1 44.41 76.68 58.04 35.96 55.47
-USC (Chen et al., 2024b) 4 48.14 80.21 61.54 37.64 58.73
-GenRead (Yu et al., 2023a) 2 42.03 74.20 56.99 34.27 53.55
-RobustRAG (Xiang et al., 2024) 11 47.80 78.09 56.29 37.08 56.53
-InstructRAG (Wei et al., 2024) 1 47.12 83.04 58.04 41.01 58.83
-Self-Route (Xu et al., 2024a) 1-2 47.46 78.80 59.09 41.01 58.06
-Astute RAG (t=1) 2 52.20 84.10 60.14 44.38 61.71
-Astute RAG (t=2) 3 53.22 84.45 61.89 44.94 62.67
-Astute RAG (t=3) 4 53.56 84.45 62.24 44.94 62.86
-Table 1: Main results on Claude under zero-shot setting, showing the accuracy of different benchmark methods vs.
-Astute RAG, along with their prediction complexity, in number of prediction API calls.
-Best scores are in bold.
+In this section, we introduce the experiment setting (Sec.
+5.1), compare the performance of Astute RAG with various baselines on diverse datasets (Sec.
+5.2), and provide in-depth analyses (Sec.
+5.3).
 5.1 Experimental Settings
 Datasets and metrics.
-We conduct experiments on the data collected in Section 2 consisting of data from NQ, TriviaQA, BioASQ, and PopQA.
-For each instance from these datasets, we provide 10 passages collected under a realistic retrieval setting: for each question in our benchmark, we query Google Search to retrieve the top 30 results and select the first 10 accessible websites.
-From each retrieved website, we extract the paragraph corresponding to the snippet provided in Google Search results as the retrieved passage..
-Most of the retrieval results contains natural noise with irrelevant or misleading information.
-We do not consider enhancements to the retrieval side, such as query rewriting, as such enhancements are typically already incorporated into commercial information retrieval systems.
-Notably, we do not select questions or annotate answers based on the retrieval results.
-This setting allows us to analyze the severity of imperfect retrieval in real-world RAG.
-It distinguishes our benchmark from previous ones that employ synthetic retrieval corruptions or that unintentionally reduce the frequency of imperfect retrieval with biased construction protocols (Chen et al., 2024a; Yang et al., 2024).
-We also evaluate our method on RGB (Chen et al., 2024a), a RAG diagnostic benchmark evaluating several crucial RAG abilities.
-Specifically, we choose the English subset of RGB focusing on noise robustness.
-The benchmark have positive and negative passage sets for each question.
-We select five negative documents per question as the context to form a worst-case scenario.
-All the data in these datasets are short-form QA.
-Following previous work (Xiang et al., 2024; Wei et al., 2024; Mallen et al., 2023), a model response is considered correct if it contains the ground-truth answer.
-To enhance evaluation reliability, we prompt LLMs to enclose the exact answer within special tokens, extracting them as the final responses.
-General Settings of LLMs and RAG.
-We conduct experiments on two advanced LLMs, including Gemini 1.5 Pro [^5] https://deepmind.google/technologies/gemini/pro/ (gemini-1.5-pro-002) and Claude 3.5 Sonnet [^6] https://www.anthropic.com/news/claude-3-5-sonnet (claude-3-5-sonnet@20240620).
-The generation temperature is set to 0 and the maximum output tokens is set to 1,024, if not specified otherwise.
-By default, the passages are presented in the prompt by reversed order.
-All experiments are under the zero-shot setting for controlled evaluation, where no demonstrations for QA or method-specific steps are provided.
+We consider datasets encompass general questions, domain-specific questions, long-tail questions, as well as both short-form and long-form formats, following prior work Xiang et al.
+(2024); Wei et al.
+(2024).
+On NQ, TriviaQA, BioASQ, and PopQA, we provide 10 passages collected with Google Search from the Web for each instance.
+For long-form QA, we use ASQA Stelmakh et al.
+(2022).
+We also evaluate on RGB (Chen et al., 2024a).
+We choose the English subset (refined version) focusing on noise robustness.
+For each instance, we select five top negative passages to form a worst-case scenario.
+Following prior work, we report the accuracy by string match.
+More details are in Appx.
+B.
+Models and General Settings.
+We conduct experiments on advanced proprietary and open-source LLMs of different scales, including Claude 3.5 Sonnet (claude-3-5-sonnet@20240620), [^5] https://www.anthropic.com/news/claude-3-5-sonnet Gemini 1.5 Pro (gemini-1.5-pro-002), [^6] https://deepmind.google/technologies/gemini/pro/ Mistral-Large (128B; version 2407), and Mistral-Nemo (12B; version 2407).
+The generation temperature is set to 0 and the maximum output tokens is set to 1,024.
+All experiments are under the zero-shot setting for controlled evaluation.
 Baselines.
-We compare Astute RAG with various RAG methods designed for enhanced robustness and representative inference strategies designed to improve response trustworthiness.
-USC (Chen et al., 2024b) is the universal self-consistency method that samples multiple LLM responses given the same context and aggregates the answers.
+We compare Astute RAG with various RAG methods designed for enhanced robustness.
+USC (Chen et al., 2024b) is a self-consistency method that samples multiple LLM responses and aggregates the answers.
 It provides a reference of naive improvements using additional API calls.
-The temperature for sampling responses in this baseline is set to 0.7.
-Genread (Yu et al., 2023a) augments retrieved passages with LLM-generated passages.
-It provide a reference of presenting passages from both internal and external knowledge in the prompt without effectively combining them.
-RobustRAG (Xiang et al., 2024) aggregates answers from each independent passage to provide certifiable robustness.
-We use the keyword aggregation variant as it is shown to be the best-performing variant on advanced LLMs.
+Genread (Yu et al., 2023a) augments retrieved passages with LLM-generated passages without explicit consolidation process.
+RobustRAG (Xiang et al., 2024) aggregates answers from independent passages to provide certifiable robustness.
+We use the best-performing keyword aggregation variant.
 InstructRAG (Wei et al., 2024) instructs the LLM to provide a rationale connecting the answer with information in passages.
-For a fair comparison, we use the instructions without training or in-context learning.
-Self-Route (Xu et al., 2024a) adaptively switches between LLMs with and without RAG.
-[^7] The original Self-Route switches between RAG and long-context LLMs, while our implementation switches between LLMs with and without RAG to better align with the problem formulation in this paper.
-This baseline provides a reference of switching between LLMs’ internal and external knowledge.
-Implementation Details of Astute RAG.
-The prompt templates for Astute RAG can be found in Appendix A.
-By default, we use 2 API calls per query, setting $t=1$ to merge the prompt for knowledge consolidation and answer finalization.
-For adaptive generation of internal knowledge, we prompt the LLM to generate no more than one passage.
-Method #API Calls NQ TriviaQA BioASQ PopQA Overall
-Gemini 1.5 Pro (002)
-No RAG 1 44.75 80.21 45.80 25.28 51.34
-RAG 1 42.71 75.97 55.24 33.71 53.65
-USC (Chen et al., 2024b) 4 46.44 76.68 58.39 37.64 56.43
-GenRead (Yu et al., 2023a) 2 45.08 77.39 54.90 34.27 54.70
-RobustRAG (Xiang et al., 2024) [^8] We observe a high refusal rate in responses of RobustRAG. 11 34.24 67.49 44.06 32.02 45.59
-InstructRAG (Wei et al., 2024) 1 46.78 80.57 54.90 34.83 56.14
-Self-Route (Xu et al., 2024a) 1-2 47.46 79.86 58.04 38.20 57.58
-Astute RAG (t=1) 2 50.17 81.63 58.04 40.45 59.21
-Astute RAG (t=2) 3 51.53 81.27 58.74 40.45 59.69
-Astute RAG (t=3) 4 48.47 80.21 60.14 42.13 59.21
-Table 2: Main results on Gemini under zero-shot setting, showing the accuracy of different benchmark methods vs.
-Astute RAG, along with their prediction complexity, in number of prediction API calls.
-Best scores are in bold.
+For a fair comparison, no training is applied.
+Self-Route (Xu et al., 2024) adaptively switches between LLMs with and without RAG.
+[^7] The original Self-Route switches between RAG and long-context LLMs, while our implementation switches between RAG and No RAG according to our problem formulation.
+It provides a reference of switching between LLMs’ internal and external knowledge.
+Implementation Details.
+The prompt templates for Astute RAG can be found in Appx.
+A.
+By default, we set $t=1$ and $\hat{m}=1$ to limit the number of additional tokens used.
+Results with larger $t$ and $\hat{m}$ are discussed in Sec.
+5.3.
+Figure 5: Performance on ASQA.
+Figure 6: Worst-case performance of Claude on RGB. Astute RAG reaches a performance close to No RAG, while other RAG systems are far behind.
 5.2 Main Results
-Table 1 and Table 2 presents the results on data with realistic retrieval augmentation for each dataset.
-By comparing RAG and No RAG, we find that retrieved passages might not always bring benefits to downstream performance – on NQ and TriviaQA, RAG performance lags behind No RAG.
-We attribute this to that the questions being covered by the LLM’s internal knowledge and the noise in retrieval results misleading the LLM.
-In contrast, on BioASQ and PopQA, which focus on domain-specific and long-tail questions, RAG significantly improves LLM performance.
-However, due to imperfect retrieval augmentation, the absolute performance still remains to be unsatisfactory.
-Among all baselines, no single method consistently outperforms others across all datasets.
-This observation highlights that these baselines are tailored to distinct settings and may not be universally applicable.
-For instance, InstructRAG is more effective on TriviaQA, achieving the best performance among all baselines with both Claude and Gemini.
-In contrast, Self-Route performs better than InstructRAG on both NQ and BioASQ.
-Moreover, RobustRAG achieves very different performance when applied to Gemini and Claude.
-Through in-depth analysis, we find that RobustRAG with Gemini exhibits a high refusal rate (refuse to answer) in responses.
-We attribute this instability to the varying method designs of the baselines, which are tailored for different scenarios, resulting in inconsistent improvement across datasets.
-Overall, InstructRAG and Self-Route demonstrates the best performance among all baselines when applied to Claude and Gemini respectively.
-We also note that increasing the number of API calls does not necessarily correlate with improved performance.
-Astute RAG consistently outperforms baselines across all datasets of different properties.
-The overall improvement compared with the best baseline is relatively 6.85% on Claude and 4.13% on Gemini, and the improvements in domain-specific questions are much higher.
-These results highlight the effectiveness of Astute RAG in overcoming imperfect retrieval augmentation.
-On Claude, adding more iteration of knowledge consolidation leads to consist improvement.
-The improvement margin becomes lower when $t$ becomes larger.
-This is because after each iteration, the remaining improvement space for knowledge consolidation becomes smaller.
-On Gemini, increasing $t$ primarily benefits BioASQ and PopQA.
-These two datasets rely more heavily on external knowledge, and iterative knowledge consolidation helps mitigate noise within this external information.
-Performance on NQ and TriviaQA does not improve further when $t$ reaches 3.
-We attribute this to the less critical role of external knowledge in these datasets.
-For setting consistency and efficiency, we set the parameter $\hat{m}$ to a smaller value, limiting the influence of internal knowledge.
+Performance under real-world retrieval.
+Tab.
+1 presents the results with real-world retrieval augmentation of various LLMs.
+We find that retrieved passages might not always bring benefits – on NQ and TriviaQA, RAG performance lags behind No RAG for advanced LLMs.
+We attribute this questions being covered by the LLM’s internal knowledge and the noise in retrieval results misleading the LLM.
+In contrast, on BioASQ and PopQA, which focus on domain-specific and ‘long-tail’ questions, RAG significantly improves the LLM performance.
+Due to imperfect retrieval augmentation, however, the absolute performance still remains to be unsatisfactory.
+Among all baselines, no single method consistently outperforms others across all datasets and LLMs.
+This observation highlights these baselines being tailored to distinct settings and not being universally applicable.
+Overall, InstructRAG and Self-Route demonstrate relatively superior performance among other alternatives.
+Astute RAG consistently outperforms baselines across all LLMs in terms of overall accuracy.
+The relative improvement compared to the best baseline is 6.85% for Claude and 4.13% for Gemini, with the improvements in domain-specific questions being much higher.
+These highlight the effectiveness of Astute RAG in overcoming imperfect retrieval augmentation and knowledge conflicts.
+Additionally, we observe consistent improvements on the open-source Mistral models.
+The results demonstrate that Astute RAG generalizes well to LLMs of smaller sizes.
+Figure 7: Performance across different retrieval precision buckets. Astute RAG is consistently better.
+Performance on long-form QA.
+We conduct additional experiments on the long-form QA dataset, ASQA.
+Fig.
+5 demonstrates that Astute RAG consistently achieves significant improvements, reinforcing its effectiveness across diverse scenarios.
+Worst-case performance on RGB.
+Fig.
+6 presents the results under the worst-case setting on RGB where all retrieved documents are negative, to demonstrate robustness.
+The performance gap between RAG and No RAG exceeds 50 points, highlighting the detrimental impact of imperfect retrieval results and emphasizing the importance of providing robust safeguards against worst-case scenarios.
+While the baseline RAG methods outperform the original RAG, they still obviously fall behind ‘No RAG’.
+Astute RAG is the only RAG method that reaches a performance close to ‘No RAG’, further supporting its effectiveness in addressing imperfect retrieval augmentation.
 5.3 Analyses
-Performance by retrieval precision.
-We compare the performance of Astute RAG and baselines across different subsets partitioned by their retrieval precision, on our collected data with Claude as the LLM.
-As shown in Figure 4, Astute RAG achieves consistently better performance than all baselines across different retrieval precision, indicating its effectiveness in improving RAG trustworthiness in broad scenarios.
+We conduct in-depth analyses using Claude following the setting of Tab. 1.
+The impact of retrieval precision.
+As shown in Fig.
+7, Astute RAG achieves consistently better performance across different retrieval precision regimes, indicating its effectiveness in improving RAG trustworthiness in broad scenarios.
 Notably, Astute RAG does not sacrifice performance gain under high retrieval quality in exchange for improvement under low retrieval quality.
-When the retrieval quality is extremely low (close to zero retrieval precision), all other RAG variants underperforms the ’No RAG’ baseline, except for the proposed Astute RAG.
-This observation aligns with the worst-case results on RGB.
-It demonstrates the difficulty in overcoming imperfect retrieval augmentation, and verify the effectiveness of Astute RAG in doing so.
-(a) Performance by Retrieval Precision
-(b) Performance by Knowledge Conflicts
-(c) Worst-case Performance on RGB
-Figure 4: Analysis results.
-(a) Performance of Claude across different buckets ranked by retrieval precision.
-(b) Performance of Claude on conflicting and consistent instances between No RAG and RAG.
-(c) Worst-case performance of Claude on RGB, where all retrieved passages are negative.
-Astute RAG reaches a performance close to No RAG, while other RAG systems are far behind.
-Effectiveness in addressing knowledge conflicts.
-We split our collected data in to three subset according to the answers from Claude, with and without RAG.
-The answers from two inference methods can be both correct, both incorrect, or conflicting with one being correct.
-These three subsets represents the three situations between internal and external knowledge.
-The results are shown in Figure 4.
-On the conflicting subset, Astute RAG successfully chooses the correct answer in approximately 80% of cases, being the most effective method in addressing knowledge conflicts.
+When the retrieval quality is extremely low (close to zero precision), all other RAG variants underperform the ’No RAG’ baseline, except for Astute RAG.
+Figure 8: Performance on conflicting and consistent instances between No RAG and RAG.
+Addressing knowledge conflicts.
+We split our collected data into three subset according to the answers with and without RAG: the answers from two can be (i) both correct, (ii) both incorrect, or (iii) conflicting with one being correct.
+The results are shown in Fig.
+8.
+On the conflicting subset, Astute RAG successfully chooses the correct answer in approximately 80% of cases, being the most effective one in addressing knowledge conflicts.
 Notably, Astute RAG even brings performance improvement on the subset where neither internal nor external knowledge alone leads to the correct answer.
-This indicates that Astute RAG can effectively combine partially-correct information from LLM-internal and external knowledge, to achieve the correct answer through collective information across them.
-Worst-case performance on RGB..
-Figure 4 presents the results under the worst-case setting on RGB where all retrieved documents are negative.
-It demonstrates the noise robustness of Astute RAG and baseline RAG methods.
-The performance gap between RAG and No RAG exceeds 50 points, highlighting the detrimental impact of imperfect retrieval results and emphasizing the importance of providing robust safeguards against worst-case scenarios.
-While the baseline RAG methods outperform the original RAG, they still obviously fall behind No RAG.
-Astute RAG is the only RAG method that reaches a performance close to No RAG under the worst-case scenario, further supporting its effectiveness in addressing imperfect retrieval augmentation.
-Qualitative study.
-In Figure 5, we present two representative examples showing the intermediate outputs of Astute RAG.
+This indicates that Astute RAG can effectively combine partially-correct information from LLM-internal and external knowledge.
+Benefits of more consolidation iteration.
+For efficiency, we employ a single iteration of knowledge consolidation in our main experiments.
+However, incorporating multiple iterations has the potential to further enhance model performance as shown in Fig.
+9.
+The magnitude of this improvement diminishes as $t$ increases, indicating that the knowledge has been better presented and less improvement space left after each iteration.
+Efficiency in tokens consumed and API calls.
+As a proxy to overall prediction cost and latency, we present the average number of tokens and API calls used per instance in Fig.
+10 and Fig.
+11.
+Astute RAG incurs only a marginal cost increase, <5%, while delivering substantial improvement, >11%, compared to the RAG baseline.
+Figure 9: Accuracy improvement when increasing $t$.
+Effectiveness of adaptive generation.
+The results in Tab.
+2 illustrate the model’s performance when varying the maximum number of passages generated.
+The design of adaptive generation has been effectively reflected, as the number of generated passages is dynamically adjusted leading to $m<\hat{m}$.
+Notably, the number of generated passages can be controlled by $\hat{m}$, and results show that the system does not generate passages excessively.
+Impact of Source-Awareness.
+To evaluate the impact of source-awareness, we conducted an ablation study where source labels (own memory and external retrieval) were removed during the consolidation and answer generation process.
+The results are shown in Tab.
+3.
+The comparison shows that providing source information (Astute RAG) leads to better performance overall compared to omitting it (Astute RAG (No Source)), particularly on NQ, PopQA, and BioASQ, suggesting that awareness of information origin aids the consolidation process.
+Accuracy of intermediate steps.
+To investigate the performance of intermediate steps, including knowledge consolidation and confidence assignment, we use LLM-as-a-judge with the instruction in Appx.
+A.
+Our experimental results show that the accuracy for knowledge consolidation is 98.2%, and for confidence assignment, it is 95.0%.
+These results demonstrate the effectiveness of the proposed framework in the intermediate stages.
+Qualitative examples.
+In Fig.
+12, we present two representative examples showing the intermediate outputs of Astute RAG.
 In the first example, LLM without RAG generates a wrong answer, while RAG returns a correct answer.
@@
 (2024).
-In the second example, LLM alone is correct, while RAG is incorrect due to the noisy retrieval results.
-Astute RAG detected the correct answer from noisy retrieved information by checking with its internal knowledge.
-Figure 5: Qualitative examples.
-Top: Astute RAG identified the error in internal knowledge (i.e., generated passage) by confirming with external sources.
-Bottom: Astute RAG detected the correct answer from noisy retrieved information by checking with its internal knowledge.
-Standard RAG does not provide an answer because the retrieved passages are too noisy.
-5 Related Work
-Retrieval augmented generation (RAG) seeks to address the inherent knowledge limitation of LLMs with passages retrieved from external sources of information such as private corpora or public knowledge bases (Guu et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022).
-Given the widespread adoption of RAG in various real-world applications, including risk-sensitive domains, the negative impact of noisy information within retrieved passages has garnered increasing attention (Cuconasu et al., 2024).
-Recent work has sought to enhance the robustness of RAG systems against noise from various perspectives, including training LLMs with noisy context (Yu et al., 2023b; Yoran et al., 2024; Pan et al., 2024; Fang et al., 2024), training small models to filter out irrelevant passages (Wang et al., 2023c; Xu et al., 2023), passage reranking (Yu et al., 2024; Glass et al., 2022), dynamic and iterative retrieval (Jiang et al., 2023; Asai et al., 2023; Yan et al., 2024), query rewriting (Ma et al., 2023), and speculative drafting (Wang et al., 2024).
-These studies focus on distinct modules or stages of RAG systems and are orthogonal to our work.
-Our work focuses on enhancing RAG robustness at the post-retrieval stage, after retrieved passages have been provided.
-On this topic, RobustRAG (Xiang et al., 2024) aggregates answers from each independent passage to provide certifiable robustness.
-InstructRAG (Wei et al., 2024) instructs the LLM to provide a rationale connecting the answer with information in passages.
-MADRA (Wang et al., 2023b) applies multi-agent debate to select helpful evidence.
-However, these works do not explicitly incorporate internal knowledge to recover from RAG failures and may therefore collapse when the majority of retrieved passages are negative.
-In terms of emphasizing internal knowledge of LLMs in RAG, recent work has explored using LLM-generated passage as context (Yu et al., 2023a; Zhang et al., 2023), adaptively switching between LLMs with and without RAG (Xu et al., 2024a; Mallen et al., 2023; Jeong et al., 2024), and combining answers from internal and external knowledge through contrastive decoding (Zhao et al., 2024; Jin et al., 2024).
-We focus on a black-box setting where no further training is required, directly addressing knowledge conflicts to combine the helpful information from both sides and achieve more reliable answers.
+In the second example, LLM is correct but RAG is incorrect due to imperfect retrieval.
+Astute RAG detected the correct answer from imperfect context leveraging internal knowledge.
+Figure 10: Efficiency in terms of tokens consumed.
 6 Conclusion
-Our paper investigates the impact of imperfect retrieval on the performance of RAG systems and identifies knowledge conflicts as a key challenge.
+We investigate the impact of imperfect retrieval on the performance of RAG systems and identify knowledge conflicts as a key challenge.
 To address this, we introduce Astute RAG, a novel approach that leverages the internal knowledge of LLMs and iteratively refines the generated responses by consolidating internal and external knowledge in a source way.
-Our empirical results demonstrate the effectiveness of Astute RAG in mitigating the negative effects of imperfect retrieval and improving the robustness of RAG systems, particularly in challenging scenarios with unreliable external sources.
-Among the limitations, Astute RAG’s effectiveness hinges on the capabilities of advanced LLMs with strong instruction-following and reasoning abilities, hence potentially more limited applicability with less sophisticated LLMs.
-As an important future direction, extending the experimental setup to include longer outputs would be important, where the challenges of imperfect retrieval and knowledge conflicts may be even more pronounced.
-Furthermore, a comprehensive analysis of the impact of various context types (Balachandran et al., 2024) would enhance the understanding of the proposed method’s effectiveness.
-Future work can also extend our method beyond LLMs and RAG, such as addressing knowledge conflicts in multimodal settings (Zhu et al., 2024).
+We demonstrate the effectiveness of Astute RAG in mitigating the negative effects of imperfect retrieval and improving the robustness of RAG, particularly in challenging scenarios with unreliable external sources.
 Acknowledgement
@@
 We would also like to thank all other colleagues from Google Cloud AI Research for their valuable feedback.
-[…]
+Limitations
+Astute RAG’s effectiveness hinges on the capabilities of advanced LLMs with strong instruction-following and reasoning abilities, hence potentially more limited applicability with less sophisticated LLMs.
+As an important future direction, extending the experimental setup to include longer inputs would be important, where the challenges of imperfect retrieval and knowledge conflicts may be even more pronounced.
 Appendix A Prompt Template for Astute RAG
@@
 Initial Context: {context_init} [Consolidated Context: {context}] # optional Question: {question} Answer:
+Prompt for Intermediate Step Evaluation **Task:** You are provided with the following: 1.
+A question.
+2.
+The correct answer.
+3.
+The input context.
+4.
+The model’s response, which contains: - Consolidated context.
+- Confidence scores for candidate answers.
+Your task is to: - Evaluate the **quality of the consolidated context** in the model’s response and provide a label: ‘<<<consolidation>>> correct <<</consolidation>>>’ or ‘<<<consolidation>>> incorrect <<</consolidation>>>’.
+This evaluation is only about whether the consolidation is correct given the input context.
+- Evaluate the **accuracy of the confidence score** (whether it aligns with the confidence of the supporting context) and provide a label: ‘<<<confidence>>> correct <<</confidence>>>’ or ‘<<<confidence>>> incorrect <<</confidence>>>’.
+The evaluation is only based on the consolidated context.
+Note that correct consolidation and confidence do not necessarily indicate the correct answer.
+Question: {query} Correct Answer: {answer} Input Context: {input} Model Response: {response} Evaluation:
 Appendix B Data Collection
@@
 Overall, our benchmark contains 1,042 short-form question-answer pairs, each paired with 10 retrieved passages.
+When collecting the passages, we retrieve the top 30 results and select the first 10 accessible websites.
+From each retrieved website, we extract the paragraph corresponding to the snippet provided in the search results as the retrieved passage.
+Retrieved results might contain natural noise with irrelevant or misleading information.
+We do not consider enhancements to the retrieval side, such as query rewriting, as such enhancements are typically already incorporated into commercial information retrieval systems.
+All of these datasets are short-form QA.
+Following previous work (Xiang et al., 2024; Wei et al., 2024; Mallen et al., 2023), a model response is considered correct if it contains the ground-truth answer.
+To enhance evaluation reliability, we prompt LLMs to enclose the exact answer within special tokens, extracting them as the final responses.
 Question-answer pairs.
@@
 We do not consider enhancements to the retrieval side, such as query rewriting, as such enhancements are typically already incorporated into commercial information retrieval systems.
+Appendix C Comparison with Answer Refinement
+In Tab.
+4, we further compare Astute RAG with Answer Refinement, where the LLM is prompted to refine its initial answer by reconsidering external context.
+Notably, this baseline performs nearly identically to Self-Route, which is expected, because both approaches rely on the model itself to determine whether to revise its initial answer based on external knowledge.
+The comparison underscores that simply enabling LLMs to self-correct does not yield significant improvements beyond existing routing strategies.
+In contrast, Astute RAG continues to outperform all baselines across datasets, reinforcing the benefit of explicitly structured consolidation mechanisms.
+Model NQ TriviaQA BioASQ PopQA Overall
+Self-Route 47.5 79.9 58.0 38.2 57.6
+Answer Refinement 47.1 79.9 58.0 38.2 57.5
+Astute RAG 50.2 81.6 58.0 40.5 59.2
+Table 4: Comparison of Astute RAG with routing and refinement-based baselines on Gemini.
+Appendix D Comparison with Context Filtering
+To isolate the effect of identifying irrelevant information separate from consolidation, we conducted an ablation study introducing two Context Filtering baselines.
+The first baseline filters only the retrieved documents, while the second filters both retrieved and generated documents prior to answer generation, both without applying the consolidation step.
+As shown in Tab.
+5, while context filtering improves performance over the basic RAG baseline, it falls short of the performance achieved by Astute RAG.
+This supports our hypothesis that consolidating diverse information, including consistent, conflicting, and relevant content, rather than merely filtering out irrelevant parts, is critical to the performance gains observed with Astute RAG.
+Model NQ TriviaQA BioASQ PopQA Overall
+RAG 42.7 76.0 55.2 33.7 53.7
+Context Filtering (retrieved) 43.7 77.0 57.0 34.3 54.8
+Context Filtering (generated and retrieved) 49.2 79.2 56.6 40.4 57.9
+Astute RAG 50.2 81.6 58.0 40.5 59.2
+Table 5: Comparison with context filtering baselines.
+Appendix E Comparison with Context Compression
+Context compression Wang et al.
+(2023b); Yoon et al.
+(2024) is also a related direction.
+We further conduct experiments comparing our method with CompAct Yoon et al.
+(2024).
+The results in Tab.
+6 and Tab.
+7 show that context compression is ineffective in handling the challenges of imperfect context and knowledge conflicts.
+Notably, it even performs worse than the No RAG and RAG baselines in this context.
+The primary reason for this underperformance lies in the limitations of context compression.
+It struggles to effectively identify incorrect information when there are conflicts in context and often removes or reduces the appearance of helpful information in the process.
+This reinforces the importance of our approach, which does not rely solely on compression but instead integrates both internal and external knowledge while handling conflicts in a more nuanced manner.
+Method NQ TriviaQA BioASQ PopQA Overall
+No RAG 47.1 82.0 50.4 29.8 54.5
+RAG 44.4 76.7 58.0 36.0 55.5
+CompAct 38.6 68.9 49.3 30.3 48.4
+Astute RAG 52.2 84.1 60.1 44.4 61.7
+Table 6: Comparison with context compression on Claude.
+Method NQ TriviaQA BioASQ PopQA Overall
+No RAG 44.8 80.2 45.8 25.3 51.3
+RAG 42.7 76.0 55.2 33.7 53.7
+CompAct 35.3 65.0 47.6 30.9 46.0
+Astute RAG 50.2 81.6 58.0 40.5 59.2
+Table 7: Comparison with context compression on Gemini.
+Appendix F Influence of passage ordering.
+We apply different ordering strategies (Alessio et al., 2024) on RAG and Astute RAG.
+As shown in Tab.
+8, we find that the improvement with Astute RAG is significantly larger than the gap between different ordering strategies.
+Moreover, the consolidation process makes Astute RAG less sensitive to it.
+Method Ordering Strategy NQ TriviaQA BioASQ PopQA Overall
+RAG Random 43.39 76.33 56.99 34.83 54.61
+Ascending 43.05 75.62 57.69 34.83 54.51
+Descending 44.41 76.68 58.04 35.96 55.47
+Ping-pong Descending Top-to-bottom 44.75 77.39 57.69 35.96 55.66
+Ping-pong Descending Bottom-to-top 44.41 75.62 58.04 35.96 55.18
+AstuteRAG Random 51.86 84.81 61.19 41.57 61.61
+Ascending 51.86 85.51 59.79 42.13 61.52
+Descending 52.20 84.10 60.14 44.38 61.71
+Ping-pong Descending Top-to-bottom 52.20 84.45 59.09 43.82 61.42
+Ping-pong Descending Bottom-to-top 51.19 85.16 61.54 43.82 62.00
+Table 8: Performance by Ordering Strategies.